# Merging RLBWTs adaptively

### v3

@@ Abstract
We show how we can merge two run-length compressed Burrows-Wheeler Transforms (RLBWTs) into a run-length compressed extended Burrows-Wheeler Transform (eBWT) in $O(r)$ space and $O((r+L)\log(m+n))$ time, where $m$ and $n$ are the lengths of the uncompressed strings, $r$ is the number of runs in the final eBWT and $L$ is the sum of the longest common prefix (LCP) values at the beginnings of those runs.
@@ 1 Introduction
As far as we know, Oliva et al. [10] were the first to study merging RLBWTs, building on Bannai, Gagie and I’s [1] work on dynamic r-indexes. Their merging algorithm uses compressed space but time at least linear in the size of the uncompressed inputs. Recently, Díaz-Domínguez et al. [5] gave an algorithm for merging BWTs that could be extended to RLBWTs and works fairly well in practice but has no good worst-case bound. In this paper we show how we can merge two RLBWTs into a run-length compressed eBWT in $O(r)$ space and $O((r+L)\log(m+n))$ time, where $m$ and $n$ are the lengths of the uncompressed strings, $r$ is the number of runs in the final eBWT and $L$ is the sum of the longest common prefix (LCP) values at the beginnings of those runs.
We describe here only how to merge the RLBWTs of two strings. Our algorithm can easily be extended to merge many RLBWTs or run-length compressed eBWTs, but we leave that to the full version of the paper. In Section 2 we review some preliminary concepts, including a technical lemma about Nishimoto and Tabei’s [9] move structures for $\Psi$ with a better bound on the construction time, which may be of independent interest. In Section 3 we present our algorithm and in Section 4 we present a preliminary analysis that yields a time bound of $O((r\log r+L)\log(m+n))$. In Section 5 we prove our technical lemma about building move structures for $\Psi$. Finally, in Section 6 we optimize our algorithm slightly to reduce the $r\log r$ to $r$, so the whole time bound becomes $O((r+L)\log(m+n))$.
@@ 2 Preliminaries
Given a character’s position in $\mathrm{BWT}_{S}$ or $\mathrm{BWT}_{T}$, we can extract its context character by character in $O(\log r)$ time plus constant time per character, if we have data structures for iteratively evaluating the $\Psi$ functions for $S$ and $T$ in $O(\log r)$ time plus constant time per iteration. Since $\Psi$ is the inverse of the LF function, Nishimoto and Tabei [9] implicitly showed how to build in $O(r)$ space such structures, which they called move structures, but they said their construction takes $O(m+n+r\log r)$ time. Brown, Gagie and Rossi [3] slightly generalized Nishimoto and Tabei’s result but did not give a construction-time bound at all. Following Brown et al.’s presentation, we show in Section 5 that, given the RLBWTs $\mathrm{BWT}_{S}$ and $\mathrm{BWT}_{T}$, we can build these structures in $O(r\log r)$ time.
Lemma 1.
@@
It follows that we can compare $\mathrm{context}_{S}(i)$ and $\mathrm{context}_{T}(j)$ in $O(\log r)$ time plus time proportional to the length of their longest common prefix.
3 Algorithm
Suppose we are to merge the RLBWTs $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}[1..n]$ of strings $S[1..m]$ and $T[1..n]$ into the run-length compressed eBWT $\mathrm{BWT}_{S,T}[1..m+n]$ of $S$ and $T$ together, and we have already built the data structures for $\Psi$ for $S$ and $T$ described in Lemma 1 so we can compare characters’ contexts. The pseudocode for our algorithm is shown in Figure 1.
@@ 3 Algorithm
Assume we have already merged $\mathrm{BWT}_{S}[1..i-1]$ and $\mathrm{BWT}_{T}[1..j-1]$ into $\mathrm{BWT}_{S,T}[1..i+j-2]$ correctly. When we first reach the while loop (lines 3–43) $i=j=1$ so this assumption is trivially true. If $i>m$ then $\mathrm{BWT}_{S,T}[i+j-1..m+n]=\mathrm{BWT}_{T}[j..n]$ (line 45) and if $j>n$ then $\mathrm{BWT}_{S,T}[i+j-1..m+n]=\mathrm{BWT}_{S}[i..m]$ (line 47), so assume $i\leq m$ and $j\leq m$ and consider only the while loop. Let $k$ and $\ell$ be the lengths of the leading runs of $\mathrm{BWT}_{S}[i..m]$ and $\mathrm{BWT}_{T}[j..n]$ respectively (lines 4 and 5).
@@
$\mathrm{BWT}_{S}[i..i+k-1]\circ\mathrm{BWT}_{T}[j..j+\ell-1]=(\mathrm{BWT}_{S}[i])^{k+j}$
@@
4 Analysis
Building the structures in Lemma 1 takes $O(r\log r)\subseteq O(r\log(m+n))$ time, where $r$ is the number of runs in $\mathrm{BWT}_{S,T}$. Since $\mathrm{BWT}_{S}$ and $\mathrm{BWT}_{T}$ both consist of at most $r$ runs, outputting $\mathrm{BWT}_{S}[i..m]$ (line 45) or $\mathrm{BWT}_{T}[j..n]$ (line 47) in run-length compressed form takes $O(r)$ time. Therefore, we can focus on the complexity of the while loop.
Lemma 2.
During each pass through he while loop, we output an entire run in $\mathrm{BWT}_{S,T}$.
Proof.
@@ 4 Analysis
$\mathrm{BWT}_{T}[j..j^{\prime}-1]=(\mathrm{BWT}_{T}[j])^{j^{\prime}-k}\,$,
the next character we output is $\mathrm{BWT}_{S}[i]\neq\mathrm{BWT}_{T}[j]$ and it starts a new run in $\mathrm{BWT}_{S,T}$. ∎
It follows from Lemma 2 that we make $O(r)$ passes through the while loop. Since $\mathrm{BWT}_{S}$ and $\mathrm{BWT}_{T}$ and $\mathrm{BWT}_{S,T}$ are run-length compressed, the only operations that can take more than constant time during a single pass through the while loop are comparisons between characters’ contexts, of which we make $O(\log(m+n))$ during each pass, so $O(r\log(m+n))$ in total.
Lemma 3.
@@
and we output $\mathrm{BWT}_{T}[j]$ during the next pass. The binary search in line 37 is symmetric. ∎
It follows from Lemma 3 that we can charge all the context comparisons to runs in $\mathrm{BWT}_{S,T}$ such that
@@
where $L$ is the sum of the LCP values at the beginnings of those runs. Combined with our previous observations, this gives us a preliminary result:
Theorem 4.
Given the run-length compressed eBWTs $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}[1..n]$, we can merge them into the run-length compressed eBWT $\mathrm{BWT}_{S,T}[1..m+n]$ using $O(r)$ space and $O((r\log r+L)\log(m+n))$ time, where $m$ and $n$ are the lengths of the uncompressed strings, $r$ is the number of runs in $\mathrm{BWT}_{S,T}$ and $L$ is the sum of the LCP values at the beginnings of those runs.
5 Proof of Lemma 1
@@
then apply Lemma 5 below to obtain a slightly larger superset $P^{\prime}$ with useful properties; and finally build a move structure with which, given $i$ and $i$’s predecessor in $P^{\prime}$, in constant time we can find $\Psi(i)$ and $\Psi(i)$’s predecessor in $P^{\prime}$. We refer readers to Nishimoto and Tabei’s [9] and Brown et al.’s [3] papers for details about constructing a move structure from $P^{\prime}$; the only significant novelty here is the improved bound on the construction time of $P^{\prime}$. Given only $i$, we can find $i$’s predecessor in $P^{\prime}$ in $O(\log r)$ time with binary search on $P^{\prime}$, then use the move structure to iterate $\psi$ and extract the context of the $i$th character in the RLBWT character by character in constant time per character; Lemma 1 follows. If the original string is over a small alphabet then we can improve the construction time in Lemmas 1 and 5, but it will not affect our final time bound for merging RLBWTs.
Lemma 5.
@@
Given the set $\{(i,\pi(i))\ :\ i\in P\}$ and an integer $d\geq 2$, in $O(|P|\log|P|)$ time we can build a set $\{(i,\pi(i))\ :\ i\in P^{\prime}\}$, where
@@
Proof.
Let $Q=\{\pi(i)\ :\ i\in P\}\cup\{n+1\}$. We build AVL trees $T_{P}$ and $T_{Q}$ storing the elements of $P$ and $Q$, respectively. For each $i\in P$ we store $\pi(i)$ as satellite data with $i$ in $T_{P}$, and we store $i$ as satellite data with $\pi(i)$ in $T_{Q}$. For each pair of consecutive elements $a$ and $b$ in $Q$, if $|[a,b)\cap P|\geq 2d$ then we store $a$ in a list $L$. This takes $O(|P|\log|P|)$ total time.
We work in rounds and maintain the relationships between $T_{P}$, $T_{Q}$ and $L$. In each round we check if $L$ is empty and, if so, we return each element $i$ in $T_{P}$ together with its satellite data $\pi(i)$; if not, we remove 1 element from $L$ and then insert 1 element into $T_{P}$, 1 element into $T_{Q}$ and up to 2 elements into $L$.
Let $P_{j}$ and $Q_{j}$ be the contents of $T_{P}$ and $T_{Q}$ after $j$ rounds, so $P_{0}=P$ and $Q_{0}=Q$. Suppose we remove $s$ from $L$ during the (j+1)(j+1)st round, meaning $s,e)\cap P_{j}\geq 2d$, where $e$ is the successor of $s$ in $Q_{j}$. We use $T_{P}$ to find the (d+1)(d+1)st largest element $p$ in $s,e)\cap P_{j}$. We use $T_{Q}$ to find $\pi^{-1}(p)$ by finding the predecessor $q$ of $p$ in $Q_{j}$ and adding pqp-q to $\pi^{-1}(q)$, which is stored as satellite data with $q$. We insert $\pi^{-1}(p)$ into $T_{P}$ with $p$ as satellite data, and insert $p$ into $T_{Q}$ with $\pi^{-1}(p)$ as satellite data. Therefore $Q_{j+1}=\{\pi(i)\ :\ i\in P_{j+1}\}\cup\{n+1\}$.
We use $T_{P}$ and $T_{Q}$ to check first whether $|[p,e)|\cap P_{j+1}|\geq 2d$; if so, we add $p$ to $L$. We then use $T_{P}$ and $T_{Q}$ to find the predecessor $s^{\prime}$ of $\pi^{-1}(p)$ in $Q_{j+1}$ and next element $e^{\prime}$ after $s^{\prime}$ in $Q_{j+1}$, and to check whether $|[s^{\prime},e^{\prime})\cap P_{j+1}|\geq 2d$; if so, we add $s^{\prime}$ to $L$. This completes the (j+1)(j+1)st round, which takes $O(\log|P_{j}|)=O(\log|Q_{j}|)=O(\log(|P|+j))$ total time.
[…]
Since the intervals in $E_{j+1}$ are disjoint and each contain at least $d$ elements of $P_{j+1}$, we have $|P_{j+1}|\geq d|E_{j+1}|\geq d(j+1)$. Since $|P_{j+1}|=|P|+j+1$ and $|P_{j+1}|\geq d(j+1)$, we have $|P|+j+1\geq d(j+1)$ and thus $j+1\leq\frac{|P|}{d-1}$ and $|P_{j+1}|=|P|+j+1\leq\frac{d|P|}{d-1}$. It follows that we return the set $\{(i,\pi(i))\ :\ i\in P^{\prime}\}$ after at most $\frac{|P|}{d-1}\in O(|P|)$ rounds and $O(|P|\log|P|)$ total time. ∎
@@ 6 Optimization
The $\log r$ in the time bound in Theorem 4 is the time to find the predecessor in a set $P^{\prime}$ with cardinality $O(r)$ of a position $i$ in an RLBWT, using binary search on $P^{\prime}$. It is multiplied by the $O(r)$ passes we make through the while loop and by the $O(\log(m+n))$ context comparisons we may make during each pass if we reach line 8, 13, 28 or 37 and perform a binary search in an RLBWT. If we do not perform such a binary search during a pass then we make only a constant number of context comparisons in that pass, and they contribute only $O(\log r)$ overhead.
To see how to speed up the binary searches in the RLBWTs, consider how we search for the smallest value $j^{\prime}$ with $\mathrm{context}_{T}(j^{\prime})\succ\mathrm{context}_{S}(i+k)$ in line 8, for example, letting $P_{T}^{\prime}$ be the set we build with Lemma 5 as part of the move structure for $\psi$ for $T$. We repeatedly choose candidate values $j^{\prime}$ in $j,j+\ell)$ and, for each one, compare $\mathrm{context}_{T}(j^{\prime})$ with $\mathrm{context}_{S}(i+k)$. If we choose the candidate values $j^{\prime}$ naïvely, then each context comparison starts with an $O(\log r)$-time predecessor search in $P_{T}^{\prime}$. If we choose $j^{\prime}\in P_{T}^{\prime}$ or already knowing the predecessor of $j^{\prime}$ in $P_{T}^{\prime}$, however, then we do not need such a search.
We first use $O(\log r)$-time binary searches in $P_{T}^{\prime}$ to find the successor of $j$ and the predecessor of $j+\ell-1$. We then use binary search to find the smallest value $p\in[j,j+\ell)\cap P_{T}^{\prime}$ with $\mathrm{context}_{T}(p)\succ\mathrm{context}_{S}(i+k)$, if there is one. Whether $p$ exists or not, we can now focus our search on interval of $j,j+\ell)$ in which all the elements have the same predecessor in $P_{T}^{\prime}$. Binary searches on $\mathrm{BWT}_{S}$ can be sped up symmetrically.
We perform two binary searches instead of one but both are in subsets of $j,j+\ell)$, so Lemma 3 still holds and we can charge all the context comparisons as before. Now, however, the time for the context comparisons charged to a run $\mathrm{BWT}_{S,T}[a..b]$ in $\mathrm{BWT}_{S,T}$ is only
@@
Theorem 6.
Given the run-length compressed eBWTs $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}[1..n]$, we can merge them into the run-length compressed eBWT $\mathrm{BWT}_{S,T}[1..m+n]$ using $O(r)$ space and $O((r+L)\log(m+n))$ time, where $m$ and $n$ are the lengths of the uncompressed strings, $r$ is the number of runs in $\mathrm{BWT}_{S,T}$ and $L$ is the sum of the LCP values at the beginnings of those runs.

### v4

@@ Abstract
We show how to merge two run-length compressed Burrows-Wheeler Transforms (RLBWTs) into a run-length compressed extended Burrows-Wheeler Transform (eBWT) in $O(r)$ space and $O((r+L)\log(m+n))$ time, where $m$ and $n$ are the lengths of the uncompressed strings, $r$ is the number of runs in the final eBWT and $L$ is the sum of its irreducible LCP values.
keywords:
Burrows-Wheeler Transform, run-length compression, RLBWT, construction, merging
@@ 1 Introduction
Merging BWTs is an established research topic [9, 10, 20] and merging RLBWTs [19, 14] is a natural extension of that. As far as we know, however, until recently all algorithms for merging RLBWTs were based on dynamic RLBWTs [18, 2] and used time at least roughly linear in the size of the uncompressed inputs, even if they used compressed space. Díaz-Domínguez et al. [8] gave an algorithm for merging BWTs that can be extended to RLBWTs and works faster when the inputs are individually repetitive but not similar to each other. Because it is based on prefix-free parsing, however, it does not have good worst-case bounds. In this paper we show how to merge two RLBWTs into a run-length compressed eBWT in $O(r)$ space and $O((r+L)\log(m+n))$ time, where $m$ and $n$ are the lengths of the uncompressed strings, $r$ is the number of runs in the final eBWT and $L$ is the sum of the longest common prefix (LCP) values at the beginnings of those runs (known as its irreducible LCP values). It is known that $L\in O((m+n)\log\delta)$ [12], where $\delta\leq r$ is a powerful measure of the eBWT’s compressibility [13].
We describe here only how to merge the RLBWTs of two strings. Our algorithm can easily be extended to merge many RLBWTs or run-length compressed eBWTs, but we leave that to the full version of the paper. In Section 2 we review some preliminary concepts. In Section 3 we present our algorithm and in Section 4 we present a first analysis that yields a time bound of $O((r\log r+L)\log(m+n))$. In Section 5 we give the proof of a technical lemma we use, whose details we refer to in Section 6 when we optimize our algorithm slightly to reduce the $r\log r$ in our time bound to $r$, so the whole bound becomes $O((r+L)\log(m+n))$.
@@ 2 Preliminaries
Given a character’s position in $\mathrm{BWT}_{S}$ or $\mathrm{BWT}_{T}$, we can extract its context character by character in $O(\log r)$ time plus constant time per character, if we have data structures for iteratively evaluating the $\Psi$ functions for $S$ and $T$ in $O(\log r)$ time plus constant time per iteration. Since $\Psi$ is the inverse of the LF function, Nishimoto and Tabei [17] implicitly showed how to build in $O(r)$ space such structures, which they called move structures; Brown et al. [6] slightly generalized their construction, and Brown [5] and Bertram et al. [3] sped it up to run in $O(r\log r)$ time.
Lemma 2.1.
@@
It follows that we can compare $\mathrm{context}_{S}(i)$ and $\mathrm{context}_{T}(j)$ in $O(\log r)$ time plus time proportional to the length of their longest common prefix. We give a proof of Lemma 2.1 in Section 5 and then explain in Section 6 how we can pay the $O(\log r)$ overhead only three times during a binary search to find the smallest value $j$ with $\mathrm{context}_{T}(j)\succ\mathrm{context}_{S}(i)$, for example, rather than paying it at every step of the search.
3 Main algorithm
Suppose we are to merge the RLBWTs $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}[1..n]$ of strings $S[1..m]$ and $T[1..n]$ into the run-length compressed eBWT $\mathrm{BWT}_{S,T}[1..m+n]$ of $S$ and $T$ together, and we have already built the data structures for $\Psi$ for $S$ and $T$ described in Lemma 2.1 so we can compare characters’ contexts. The pseudocode for our algorithm is shown in Figure 1.
@@ 3 Main algorithm
Assume we have already merged $\mathrm{BWT}_{S}[1..i-1]$ and $\mathrm{BWT}_{T}[1..j-1]$ into $\mathrm{BWT}_{S,T}[1..i+j-2]$ correctly. When we first reach the while loop (lines 3–43 in Figure 1) $i=j=1$ so this assumption is trivially true. If $i>m$ then $\mathrm{BWT}_{S,T}[i+j-1..m+n]=\mathrm{BWT}_{T}[j..n]$ (line 45) and if $j>n$ then $\mathrm{BWT}_{S,T}[i+j-1..m+n]=\mathrm{BWT}_{S}[i..m]$ (line 47), so assume $i\leq m$ and $j\leq n$ and consider only the while loop. Let $k$ and $\ell$ be the lengths of the leading runs of $\mathrm{BWT}_{S}[i..m]$ and $\mathrm{BWT}_{T}[j..n]$ respectively (lines 4 and 5).
@@
$\mathrm{BWT}_{S}[i..i+k-1]\circ\mathrm{BWT}_{T}[j..j+\ell-1]=(\mathrm{BWT}_{S}[i])^{k+\ell}$
@@
4 First analysis
Building the structures in Lemma 2.1 takes $O(r\log r)\subseteq O(r\log(m+n))$ time, where $r$ is the number of runs in $\mathrm{BWT}_{S,T}$. Since $\mathrm{BWT}_{S}$ and $\mathrm{BWT}_{T}$ both consist of at most $r$ runs, outputting $\mathrm{BWT}_{S}[i..m]$ (line 45) or $\mathrm{BWT}_{T}[j..n]$ (line 47) in run-length compressed form takes $O(r)$ time. Therefore, we can focus on the complexity of the while loop.
Lemma 4.1.
During each pass through the while loop, we output an entire run in $\mathrm{BWT}_{S,T}$.
Proof 4.2.
@@ 4 First analysis
$\mathrm{BWT}_{T}[j..j^{\prime}-1]=(\mathrm{BWT}_{T}[j])^{j^{\prime}-j}\,$,
the next character we output is $\mathrm{BWT}_{S}[i]\neq\mathrm{BWT}_{T}[j]$ and it starts a new run in $\mathrm{BWT}_{S,T}$.
It follows from Lemma 4.1 that we make $O(r)$ passes through the while loop. Since $\mathrm{BWT}_{S}$ and $\mathrm{BWT}_{T}$ and $\mathrm{BWT}_{S,T}$ are run-length compressed, the only operations that can take more than constant time during a single pass through the while loop are comparisons between characters’ contexts, of which we make $O(\log(m+n))$ during each pass, so $O(r\log(m+n))$ in total.
Lemma 4.3.
@@
and we output $\mathrm{BWT}_{T}[j]$ during the next pass. The binary search in line 37 is symmetric.
It follows from Lemma 4.3 that we can charge all the context comparisons to runs in $\mathrm{BWT}_{S,T}$ such that
@@
where $L$ is the sum of its irreducible LCP values. Combined with our previous observations, this gives us a preliminary result:
Theorem 4.5.
Given the run-length compressed eBWTs $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}[1..n]$, we can merge them into the run-length compressed eBWT $\mathrm{BWT}_{S,T}[1..m+n]$ using $O(r)$ space and $O((r\log r+L)\log(m+n))$ time, where $m$ and $n$ are the lengths of the uncompressed strings, $r$ is the number of runs in $\mathrm{BWT}_{S,T}$ and $L$ is the sum of its irreducible LCP values.
5 Proof of Lemma 2.1
The $\log r$ in Theorem 4.5’s time bound comes from paying the $O(\log r)$ overhead in Lemma 2.1 at every step of each binary search. In Section 6 we will explain how to avoid that and pay the overhead only three times during a binary search, but first it helps to examine the proof of Lemma 2.1.
@@
then apply Lemma 5.1 below to obtain a slightly larger superset $P^{\prime}$ with useful properties; and finally build a move structure with which, given $i$ and $i$’s predecessor in $P^{\prime}$, in constant time we can find $\Psi(i)$ and $\Psi(i)$’s predecessor in $P^{\prime}$. Given only $i$, we can find $i$’s predecessor in $P^{\prime}$ in $O(\log r)$, then use the move structure to iterate $\psi$ and extract the context of the $i$th character in the RLBWT character by character in constant time per character; Lemma 2.1 follows.
We refer readers to Nishimoto and Tabei’s [17], Brown et al.’s [6] and Bertram et al.’s [3] papers and Brown’s [5] thesis for more details on building move structures. For the sake of completeness, we include a proof of Lemma 5.1 in the appendix, mostly following Brown et al.’s and Bertram et al.’s presentation but with a potential-function argument by Alhadi [1].
Lemma 5.1.
@@
Given the set $\{(i,\pi(i))\ :\ i\in P\}$ and an integer $d\geq 2$, we can build a set $\{(i,\pi(i))\ :\ i\in P^{\prime}\}$, where
@@
This takes $O(|P|\log|P|)$ time in general but $O\left(\frac{|P|\log|P|}{d}\right)$ time when $n$ is polynomial in $|P|$.
@@ 6 Optimization
The $O(\log r)$ overhead in Lemma 2.1 comes from a predecessor query on a set of size $O(r)$, as described in Section 5. For our algorithm, that is multiplied by the $O(r)$ passes we make through the while loop and by the $O(\log(m+n))$ context comparisons we may make during each pass if we reach line 8, 13, 28 or 37 and perform a binary search in an RLBWT. If we do not perform such a binary search during a pass then we make only a constant number of context comparisons in that pass and they contribute only $O(\log r)\subseteq O(\log(m+n))$ overhead, so we need not worry about this case.
To see how to speed up the binary searches in the RLBWTs, consider how we search for the smallest value $j^{\prime}$ with $\mathrm{context}_{T}(j^{\prime})\succ\mathrm{context}_{S}(i+k)$ in line 8, for example. We repeatedly choose candidate values $j^{\prime}$ in $j,j+\ell)$ and, for each one, compare $\mathrm{context}_{T}(j^{\prime})$ with $\mathrm{context}_{S}(i+k)$. If we choose the candidate values $j^{\prime}$ naïvely, then each context comparison starts with a predecessor query on $P_{T}^{\prime}$, where $P_{T}^{\prime}$ is the set we build with Lemma 5.1 as part of building a move structure for $\Psi$ on $T$. If we choose $j^{\prime}\in P_{T}^{\prime}$ or already knowing the predecessor of $j^{\prime}$ in $P_{T}^{\prime}$, however, then we do not need such a predecessor query and we do not pay the $O(\log n)$ overhead for that step of the binary search.
In an optimized binary search for $j^{\prime}$, we first pay the $O(\log r)$ overhead once to be able to extract $\mathrm{context}_{S}(i+k)$. We then pay $O(\log r)$ overheads twice more to find the successor of $j$ and the predecessor of $j+\ell-1$ in $P_{T}^{\prime}$. We then use binary search to find the smallest value $p\in[j,j+\ell)\cap P_{T}^{\prime}$ with $\mathrm{context}_{T}(p)\succ\mathrm{context}_{S}(i+k)$, if there is one. Since $p\in[j,j+\ell)\cap P_{T}^{\prime}\subseteq P_{T}^{\prime}$, we do not pay the $O(\log n)$ overhead for these steps. Whether $p$ exists or not, we can now focus our search on interval of $j,j+\ell)$ in which all the elements have the same predecessor in $P_{T}^{\prime}$, so we do not pay the $O(\log n)$ overhead for these steps either. Binary searches on $\mathrm{BWT}_{S}$ can be sped up symmetrically.
Lemma 4.3 still holds and we can charge all the context comparisons as before. Now, however, the time for the context comparisons charged to a run $\mathrm{BWT}_{S,T}[a..b]$ in $\mathrm{BWT}_{S,T}$ is only
@@
Theorem 6.1.
Given the run-length compressed eBWTs $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}[1..n]$, we can merge them into the run-length compressed eBWT $\mathrm{BWT}_{S,T}[1..m+n]$ using $O(r)$ space and $O((r+L)\log(m+n))$ time, where $m$ and $n$ are the lengths of the uncompressed strings, $r$ is the number of runs in $\mathrm{BWT}_{S,T}$ and $L$ is the sum of its irreducible LCP values.
Appendix A Proof of Lemma 5.1
Lemma A.0.
Let $\pi$ be a permutation on $\{1,\ldots,n\}$ and
$P=\{1\}\cup\{i\ :\ 1<i\leq n,\pi(i)\neq\pi(i-1)+1\}\,.$
Given the set $\{(i,\pi(i))\ :\ i\in P\}$ and an integer $d\geq 2$, we can build a set $\{(i,\pi(i))\ :\ i\in P^{\prime}\}$, where
  • $P\subseteq P^{\prime}\subseteq\{1,\ldots,n\}$,
  • $|P^{\prime}|\leq\frac{d|P|}{d-1}$.
This takes $O(|P|\log|P|)$ time in general but $O\left(\frac{|P|\log|P|}{d}\right)$ time when $n$ is polynomial in $|P|$.
Proof A.1.
Let $Q=\{\pi(i)\ :\ i\in P\}\cup\{n+1\}$. We build AVL trees $T_{P}$ and $T_{Q}$ storing the elements of $P$ and $Q$, respectively. For each $i\in P$ we store $\pi(i)$ as satellite data with $i$ in $T_{P}$, and we store $i$ as satellite data with $\pi(i)$ in $T_{Q}$. For each pair of consecutive elements $a$ and $b$ in $Q$, if $|[a,b)\cap P|\geq 2d$ then we store the pair $(a,b)$ in a list $L$. The bottleneck is sorting $P$ and $Q$, which takes $O(|P|\log|P|)$ time in general but $O(|P|)$ time when $n$ is polynomial in $|P|$.
We work in rounds and maintain the relationships between $T_{P}$, $T_{Q}$ and $L$. In each round we check if $L$ is empty and, if so, we return each element $i$ in $T_{P}$ together with its satellite data $\pi(i)$ and then stop; if not, we remove 1 element from $L$ and then insert 1 element into $T_{P}$, 1 element into $T_{Q}$ and up to 2 elements into $L$.
Let $P_{j}$ and $Q_{j}$ be the contents of $T_{P}$ and $T_{Q}$ after $j$ rounds, so $P_{0}=P$ and $Q_{0}=Q$. Suppose we remove $(s,e)$ from $L$ during the (j+1)(j+1)st round, meaning $s,e)\cap P_{j}\geq 2d$. We use $T_{P}$ to find the (d+1)(d+1)st largest element $p$ in $s,e)\cap P_{j}$. We use $T_{Q}$ to find $\pi^{-1}(p)$ by finding the predecessor $q$ of $p$ in $Q_{j}$ and adding pqp-q to $\pi^{-1}(q)$, which is stored as satellite data with $q$. We insert $\pi^{-1}(p)$ into $T_{P}$ with $p$ as satellite data, and insert $p$ into $T_{Q}$ with $\pi^{-1}(p)$ as satellite data. Therefore $Q_{j+1}=\{\pi(i)\ :\ i\in P_{j+1}\}\cup\{n+1\}$.
To see why $\pi^{-1}(p)=\pi^{-1}(q)+p-q$ and suppose $\pi(x)=y$. If $x\in P_{j}$ then $y\in Q_{j}$, so if $y\not\in Q_{j}$ then $x\not\in P_{j}$. Therefore, if $y\not\in Q_{j}$ then $\pi(x)=\pi(x-1)+1$ and so $y-1=\pi(x-1)$; applying $\pi^{-1}$ to both sides we have $\pi^{-1}(y-1)=x-1$ and so $\pi^{-1}(y)=\pi^{-1}(y-1)+1$. Since $q$ is $p$’s predecessor in $Q_{j}$, it follows that $\pi^{-1}(p)=\pi^{-1}(q)+p-q$.
We use $T_{P}$ and $T_{Q}$ to check first whether $|[p,e)\cap P_{j+1}|\geq 2d$; if so, we add $(p,e)$ to $L$. We then use $T_{P}$ and $T_{Q}$ to find the predecessor $s^{\prime}$ of $\pi^{-1}(p)$ in $Q_{j+1}$ and next element $e^{\prime}$ after $s^{\prime}$ in $Q_{j+1}$, and to check whether $|[s^{\prime},e^{\prime})\cap P_{j+1}|=2d$; if so, we add $(s^{\prime},e^{\prime})$ to $L$. Notice that if $|[s^{\prime},e^{\prime})\cap P_{j+1}|>2d$ then $|[s^{\prime},e^{\prime})\cap P_{j}|\geq 2d$ so $(s^{\prime},e^{\prime})$ was already in $L$. This completes the (j+1)(j+1)st round, which takes $O(\log|P_{j}|)$ total time.
To bound the number of rounds we use, let
[…]
$2d\leq|[s,e)\cap P_{j}|=|[s,p)\cap P_{j}|+|[p,e)\cap P_{j}|=d+|[p,e)\cap P_{j}|\,$,
we have
[…]
It follows that $f(j+1)\leq f(j)-(d-1)$.
Since $f(0)\leq|P|$ and we stop if $f(j+1)=0$ and $|P_{j+1}|\leq|P_{j}|+1$, we use at most $\frac{|P|}{d-1}$ rounds and return $P^{\prime}$ with
$|P^{\prime}|\leq|P|+\frac{|P|}{d-1}=\frac{d|P|}{d-1}\,.$
Since the (j+1)(j+1)st round takes $O(\log|P_{j}|)=O(\log|P|)$ time for all $j$, over all the rounds we use $O\left(\frac{|P|\log|P|}{d}\right)$ total time.